January 1 of the current year, Phobos Company acquired 10,000 shares of Investment in equity designated as a Fair Value through Other Comprehensive Income of Deimos Company at P 400,000 plus brokerage expenses of P 20,000. On March 1 of the current year, Deimos Company ordinary share was split on a 5 for 2 basis. On October 1, Deimos Company made a Special assessment of P 3.20 per share on all ordinary shareholder. Phobos Company accordingly paid the assessment. The fair value of December 31 amounted to P 30 per share:


1.The necessary entries on January 1, will include a

a. Debit to financial asset at FVTOCI, P 400,000

b. Debit to financial asset to FVTOCI, P 420,000

c. Credit to cash, P 400,000

d. No journal entry

2. The necessary entries on December 31, will include

a. Debit to financial asset at FVTOCI, P 400,000

b. Debit to financial asset at FVTOCI, P 300,000

c. Credit to Unrealized gain, P 250,000

d. Credit to Unrealized gain, P 140,000



1
Expert's answer
2021-06-08T12:19:58-0400


1.

Comprehensive income +brokage expenses

"400000+20000"

"=420000"


2. The necessary entries on December 31, will include


The unrealized gain = Fair value at the end of the year – Carrying amount

Fair value at the end of the year = Total shares outstanding x fair value per share

Fair value at the end of the year = 25,000 "\times" 30 = 750,000

Carrying amount = Acquisition value + brokerage fees + (assessment per share "\times" No. of outstanding shares)

Carrying amount = 400,000 + 20,000 + (3.20 "\times" 25,000)

Carrying amount = 420,000 + 80,000 = 500,000

The unrealized gain to be presented in the other comprehensive income for the current year = 750,000 – 500,000 = 250,000

  Credit to Unrealized gain, P 250,000